OpenStudy (anonymous):

4x(x+1) = 7 (solve quadratic equation)

hartnn (hartnn):

do u know quadratic formula?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but don't I need to foil this first?

hartnn (hartnn):

yes 4x (x+1) = 4x^2 + 4x

OpenStudy (anonymous):

minus 7 too right?

hartnn (hartnn):

so your quadratic equation in the form ax^2+bx+c = 0 will be 4x^2 +4x-7 = 0 yes find a,b,c values

OpenStudy (anonymous):

I get all that but I need the answer in square roots. I can't write decimals on the test.. that I don't have a problem with

hartnn (hartnn):

how are you getting decimals ? did u calculate b^2-4ac = ... ?

OpenStudy (anonymous):

I got square root of 44

hartnn (hartnn):

no... what values of a,b,c, u got ?

OpenStudy (anonymous):